Write a quadratic equation in the variable x having the given numbers
-7,3
The equation is _________________=0

Answer by checkley77(12844)

-7,3
The equation is [x^2+4x-21]=0
Factored:
(x+7)(x-3)=0
x+7=0
x=-7 ans.
x-3=0
x=3 ans.